solve by the addition method:
(x/4) - (y-3) = -4 and (x/8) + (y/6) = 0

Answer by mananth(16946)

(x/4) - (y-3) = -4
LCM = 4
x-4(y-3) / 4 =-4
x-4y+12 / 4 =-4
x-4y+12=-16
x-4y=-16-12
x-4y=-28...................1
(x/8) + (y/6) = 0
LCM 24
3x+4y /24 =0
3x+4y=0....................2
Add equation1 & 2
x-4y+3x+4y=-28
4x=-28
divide by4
4x/4=-28/4
x=-7
..
plug the value of x in equation 1
-7-4y=-28
-4y=-28+7
-4y=-21
divide by -4
y=21/4